# Multi-modal Bayesian Neural Network Surrogates with Conjugate Last-Layer Estimation

### v1

@@ 3.2 Conjugate Last Layers with Missing Data
The approach outlined in the previous section only works when the full conditional distribution, $p(\bm{W},\bm{b},\bm{\Sigma}|\bm{Z},\bm{Y})$, has a closed form. However, if some of the observations, $\bm{y}_{i}$, are only partially observed, no such closed-form distribution exists in general. For the motivating example of BO, we expect missing data to be a common phenomenon in practice. Missing observations result in partially observed response vectors in the joint model, where any $\bm{x}$ where not all modalities have been observed will have a corresponding joint $\bm{y}$ with missing values.
@@ 4.1 Simulated Data
  • Wind: The wind dataset is built from wind time series extracted from the ERA-5 model dataset [11] and measurements from Argonne National Laboratory tower measurements accessible at https://www.anl.gov/evs/atmos comprising hourly wind speed and direction. For these data the input is 1-dimensional (time) and the output of each modality is scalar. Measured wind speed at 10 meters is the main modality, and measurements at other altitudes, and modeled wind speed and direction are the auxiliary modalities.
@@
  • Time Series (1d): This dataset is the same synthetic dataset as above, but considering one scalar summary of the time series as the main quantity of interest and the PCA-reduced time series, and other summaries as auxilliary modalities.
@@
High-dimensional data presents a challenge for our proposed models, due to the presence of the Inverse-Wishart covariance matrix in (28). For an output with dimension $k$, this parameter will have dimension $k\times k$, and it is difficult to calculate the conjugate Wishart parameters and sample this matrix without numerical error for large $p$. Therefore, when modeling high-dimensional data, we first reduce the dimension through PCA. We choose PCA due to its simplicity, speed and effectiveness, in terms of limited need for tuning, its ease of decomposition and reconstruction of the data, and the resulting reduction in data size in our test datasets. Other dimension reduction methods could also be used. For a high-dimensional modality $\bm{y}^{(m)}$, we perform a PCA analysis on all observations of $\bm{y}^{(m)}$ in the training data. Then, retaining the first $c$ largest components to explain at least 95% of the variance in the data, we replace each observation $\bm{y}^{(m)}$ with a $c$-dimensional vector $\bm{y}^{(m)}_{PCA}$ of the coefficients of these $c$ components. The vectors $\bm{y}^{(m)}_{PCA}$ are used in model training. For model evaluation, we project the validation points of that modality onto the same space of principal components, keeping the same $c$ coefficients, and compare to the model’s predictions in the reduced space. Dimension reduction was necessary for the Time Series, Time Series (1d), and Wind (daily) datasets. The Time Series dataset’s high-dimensional time series modality was reduced from length 200 to length 7. In the Wind (daily) dataset, each modality was originally represented by a vector of length 24 (hourly observations for one day) and were reduced to between 3 and 10 dimensions depending on the modality.
@@ 4.2 Numerical Experiment Framework
For each dataset, we fit a joint model and a layered model to the full multi-modal data. We also fit a single BNN of the form in (2)-(4) on just the main modality data with the goal of benchmarking our multi-modal surrogates against the most similar uni-modal BNN surrogate available. Each model was fit 20 times using different initial seeds to account for the stochastic nature of the estimation procedure. Each neural network comprising the models had two hidden layers of 256 nodes each. We qualitatively saw negligible change in performance for deeper or wider networks, thus, selected 2 layers of 256 nodes as the architecture for these experiments. Each model was randomly initialized and trained until there was no statistically significant slope in the loss function over each epoch. After training, we drew 500 samples of the parameters from the trained variational posterior to use to evaluate the model’s predictive performance.
@@ 4.3 Numerical Experiment Results
Figure 2: Example of uni-modal (top) and Layered multi-modal (bottom) models fit to the same wind dataset. Notice the visible reduction in average prediction error (visualized as the difference between the blue and green lines) in the layered multi-modal model where there are no observations of the main quantity of interest, relative to the uni-modal model. Additionally, the posterior prediction intervals (visualized as the light green bands) more often captures the true function.
The multi-modal surrogate models show promising improvement on some datasets over the uni-modal model. For both Branin and Paciorek datasets, the layered model greatly reduces the bias of the predictions compared to the uni-modal model, and the joint model slightly reduces bias (Figure 3). The difference is particularly noticeable for test points outside of the convex hull of the main modality training data. The multi-modal models also perform at least as well as the uni-modal model in terms of standardized error, which we expect to be closer to one if model uncertainty is well-calibrated (Figure 4). Interestingly, both multi-modal models perform very similarly in all Paciorek datasets, regardless of the correlation of the auxiliary modalities. The layered model shows similar results on the wind dataset, while the joint multi-modal model has higher prediction bias (but better-calibrated errors) than the uni-modal model for out-of-sample test points. Figure 2 shows one trained instance of the layered multi-modal model improving out-of-sample predictions on the wind dataset due to the presence of multi-modal data in the gaps between the main modality training points. The layered model improves prediction bias and standardized error in the Wind (Daily) dataset, although both measures of error are overall higher than in other datasets. On both Time Series datasets, the layered multi-modal model performs worse in terms of average bias than both the uni-modal and joint multi-modal models, which perform comparably to each other. On the Time Series (1d) dataset, the layered model appears to have standardized errors closer to one for the In Sample and In Hull test points, but farther from one for the Out of Hull test points.
@@ 5 Conclusion
In this paper, we have developed two novel Bayesian neural network-based multi-modal surrogate models. These models build upon existing work in multi-modal learning and multi-fidelity surrogate modeling. We developed a novel family of approximate distributions for variational Bayes estimation of these models’ posterior distributions that leverages known and closed-form expressible posterior distributions in the last layer. We developed a novel technique to complete this estimation procedure in the presence of missing data by drawing the missing values from distributions that approximate the posterior predictive distribution of the missing data. Without this step, the missing data would make the last-layer posterior distributions have no known closed form. We demonstrate through simulation that these models show improved estimation for the main modality in some datasets when additional samples of auxiliary modalities are present. The canonical correlation between the auxiliary modality data, collectively, and the main modality or quantity of interest appears to be indicative to determine a priori whether these multi-modal models may lead to improved predictions compared to uni-modal alternatives. While multi-fidelity surrogate models are often hierarchical, these multi-modal models consider alternate data sources equally and enable utilizing them in combination. A method that measures non-linear relationships in order to determine a priori the performance of these multi-modal models is left for future work.
@@ Acknowledgments
This work was authored by the National Renewable Energy Laboratory, operated by Alliance for Sustainable Energy, LLC, for the U.S. Department of Energy (DOE) under Contract No. DE-AC36-08GO28308. Funding was provided by the Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program through the FASTMath Institute. The views expressed in this article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes.
This research was performed using computational resources sponsored by the Department of Energy’s Office of Energy Efficiency and Renewable Energy and located at the National Renewable Energy Laboratory.
@@ Appendix A Simulated Datasets
For the “Time Series” dataset, we use a PCA-decomposed representation of the full-length $N$ noisy time series as the main modality, and all three scalar values as the auxiliary modalities. For “Time Series (1D)”, we instead use Total Distance as the main modality and the PCA-decomposed time series and the other two scalar values as auxiliary modalities. We have one option with a scalar quantity of interest to better replicate downstream tasks such as optimization. Of all the scalar modalities, total distance is most affected by all of the input values. In both datasets, we observe the main modality at the points $(\log\alpha,\log\gamma,\mathrm{atanh}(\delta))\in\{-1.5,-1,-0.5,0.5,1,1.5\}^{3}$ and the auxiliary modalities at the points $(\log\alpha,\log\gamma,\mathrm{artanh}(\delta))\in\{-2,-1.5,-1,-0.5,0,0.5,1,1.5,2\}^{3}$. The number of PCA components is chosen to explain 95% of the variability in the observed time series. Figure 5 shows an example of the original noisy time series data along with the cumulative sum of the seven principal components that account for 95% of the variation in the training data.

### v2

@@ 3.2 Conjugate Last Layers with Missing Data
The approach outlined in the previous section only works when the full conditional distribution, $p(\bm{W},\bm{b},\bm{\Sigma}|\bm{Z},\bm{Y})$, has a closed form. However, if some of the target variables, $\bm{y}_{i}$, are only partially observed, no such closed-form distribution exists in general. The main case when missing data arises is in the joint model. Due to the concatenation of target values $\bm{y},\bm{y}^{(1)},\dots,\bm{y}^{(M)}$ into $\bm{y}^{\prime}$, any $\bm{x}$ values where only some of the modalities have been observed will have a corresponding $\bm{y}^{\prime}$ with missing values. For the motivating example of BO, we expect missing data to be a common phenomenon in practice due to not all data sources being observed at the same input values as the algorithm progresses. Missing observations result in partially observed response vectors in the joint model, where any $\bm{x}$ where not all modalities have been observed will have a corresponding joint $\bm{y}$ with missing values.
@@ 4.1 Simulated Data
  • Wind: The wind dataset is built from wind time series extracted from the ERA-5 model dataset [7] and measurements from Argonne National Laboratory tower measurements accessible at https://www.anl.gov/evs/atmos comprising hourly wind speed and direction. For these data the input is 1-dimensional (time) and the output of each modality is scalar. Measured wind speed at 10 meters is the main modality, and measurements at other altitudes, and modeled wind speed and direction are the auxiliary modalities. This illustrates a missing data imputation procedure where missing observations are infilled from regularly available model data and other observations. In practice, measurements are frequently missing, linear interpolation between missing data without additional sources of information is commonly used in these situations. However, this is poorly informative and fully deterministic.
@@
  • Time Series (1d): This dataset is the same synthetic dataset as above, but considering one scalar summary of the time series as the main quantity of interest and the PCA-reduced time series, and other summaries as auxiliary modalities.
@@
High-dimensional data presents a challenge for our proposed models, due to the presence of the Inverse-Wishart covariance matrix in (29). For an output with dimension $k$, this parameter will have dimension $k\times k$, and it is difficult to calculate the conjugate Wishart parameters and sample this matrix without numerical error for large $k$. Therefore, when modeling high-dimensional data, we first reduce the dimension through PCA. We choose PCA due to its simplicity, speed and effectiveness, in terms of limited need for tuning, its ease of decomposition and reconstruction of the data, and the resulting reduction in data size in our test datasets. Other dimension reduction methods could also be used. For a high-dimensional modality $\bm{y}^{(m)}$, we perform a PCA analysis on all observations of $\bm{y}^{(m)}$ in the training data. Then, retaining the first $c$ largest components to explain at least 95% of the variance in the data, we replace each observation $\bm{y}^{(m)}$ with a $c$-dimensional vector $\bm{y}^{(m)}_{PCA}$ of the coefficients of these $c$ components. The vectors $\bm{y}^{(m)}_{PCA}$ are used in model training. For model evaluation, we project the validation points of that modality onto the same space of principal components, keeping the same $c$ coefficients, and compare to the model’s predictions in the reduced space. Dimension reduction was necessary for the target variables in the Time Series, Time Series (1d), and Wind (daily) datasets. The Time Series dataset’s high-dimensional time series modality was reduced from length 200 to length 7. For the “Wind (Daily)” dataset, while each target variable only has dimension 24, there are 17 total variables of this size used, for a total size of 408 dimensions across all target variables. We reduce the dimension of each target variable using PCA to between 3 and 10, for a total of 94 dimensions across all 17 modalities.
@@ 4.2 Numerical Experiment Framework
For each dataset, we fit a joint model (Joint; Section 2.1) and a layered model (Layered; Section 2.2) to the full multi-modal data. We also fit a single BNN of the form in (2)-(4) on just the main modality data (Single). Comparing each model to the Single BNN allows for benchmarking our multi-modal surrogates against the most similar uni-modal BNN surrogate available. Each model was fit 20 times using different initial seeds to account for the stochastic nature of the estimation procedure. When constructing the training data for the joint model, we assumed the target variable defined in (5), $\bm{y}^{\prime}$, had missing values wherever the auxiliary modalities were observed but the main modality was not. These missing values were handled as described in 3.2. Each neural network comprising the models had two hidden layers of 256 nodes each. We qualitatively saw negligible change in performance for deeper or wider networks, thus, selected 2 layers of 256 nodes as the architecture for these experiments. Each model was randomly initialized and trained until there was no statistically significant slope in the loss function over each epoch. After training, we drew 500 samples of the parameters from the trained variational posterior to use to evaluate the model’s predictive performance.
@@ 4.3 Numerical Experiment Results
Figure 2: Example of uni-modal (top) and Layered multi-modal (bottom) models fit to the same wind dataset. Notice the visible reduction in average prediction error (visualized as the difference between the blue and green lines) in the layered multi-modal model where there are no observations of the main quantity of interest, relative to the uni-modal model. Because auxiliary data was used in training in the gaps between main modality training data (see Figure 3), the mean prediction is able to more closely follow the true function in those areas, the posterior prediction intervals (visualized as the light green bands) more often captures the true function.
Figure 3: Auxiliary data that was used to train the layered multi-modal model on the Wind dataset. The auxiliary data is measured both where the main modality is measured and also where the main modality is not. Therefore the auxiliary network is able to assist the main network in making predictions where there is no main modality training data.
The multi-modal surrogate models show promising improvement on some datasets over the uni-modal model. For both Branin and Paciorek datasets, the layered model greatly reduces the bias of the predictions compared to the uni-modal model, and the joint model slightly reduces bias (Figure 4). The difference is particularly noticeable for test points outside of the convex hull of the main modality training data. The multi-modal models also perform at least as well as the uni-modal model in terms of standardized error, which we expect to be closer to one if model uncertainty is well-calibrated (Figure 5). Interestingly, both multi-modal models perform very similarly in all Paciorek datasets, regardless of the correlation of the auxiliary modalities. The layered model shows similar results on the wind dataset, while the joint multi-modal model has higher prediction bias (but better-calibrated errors) than the uni-modal model for out-of-sample test points. Figure 2 shows one trained instance of the layered multi-modal model improving out-of-sample predictions on the wind dataset due to the presence of multi-modal data in the gaps between the main modality training points (Figure 3). In this fitted example, the auxiliary training data affects both the mean prediction and the prediction uncertainty. The mean prediction follows the true mean more closely where there is no main modality training data. In areas where the auxiliary training data is less variable ($t=3$ to $t=4$), the prediction uncertainty is decreased, while in areas where the auxiliary training data is more variable ($t=5$ to $t=6$), the uncertainty is increased. The layered model improves prediction bias and standardized error in the Wind (Daily) dataset, although both measures of error are overall higher than in other datasets. On both Time Series datasets, the layered multi-modal model performs worse in terms of average bias than both the uni-modal and joint multi-modal models, which perform comparably to each other. On the Time Series (1d) dataset, the layered model appears to have standardized errors closer to one for the In Sample and In Hull test points, but farther from one for the Out of Hull test points.
@@
5 Application Datasets
We apply these models to two real-world datasets: a dataset of building heating and cooling energy demand [[, ENB;]]tsanas_accurate_2012 and a dataset of heavy metal concentrations in the Jura region of Switzerland [[, Jura;]]goovaerts_geostatistics_1997.
5.1 ENB Dataset
The ENB dataset contains heating and cooling demand for 768 generated buildings as simulated by Autodesk Ecotect Analysis. Ecotect is a building performance analysis tool widely used by architects, engineers, and sustainability consultants.  [^3] Autodesk Ecotect Analysis has been discontinued as of March 20, 2015. (https://www.autodesk.com/support/technical/article/caas/sfdcarticles/sfdcarticles/Ecotect-Analysis-Discontinuation-FAQ.html). The buildings were generated with the same overall volume, location, interior characteristics and building materials, but with differing layout configurations, glazing areas and distributions, and cardinal orientations. The input variables recorded were relative compactness, surface area, wall area, roof area, overall height, orientation, glazing area, and glazing area distribution. These building configurations result in 586 unique heating loads and 636 unique cooling loads.
[29] use these data to demonstrate the applicability of random forest regression models over linear regression models for predicting building energy use, finding that random forests outperform linear regression due to the complex input-output relationships in the data. While in the original analysis, both variables are always measured, and there is no difference in computational difficulty between the heating and cooling simulation, one can suppose a different setting where this is not the case. For example, heating demand may be directly measured on buildings in the winter, before the study is expanded and cooling demand is measured on those and more buildings in the summer. In this setting, our multi-modal surrogate models may be trained on this dataset to predict heating demand both on new buildings and on buildings for which only cooling demand was measured. By comparison, waiting until winter to measure heating demand for the new buildings is undesirable. In our analysis, we treat heating demand as the main modality and cooling demand as an auxiliary modality. Because both heating and cooling demand are strictly positive in these data, we log-transform each variable before fitting models. Results are shown for the $\log(\bm{y})$ rather than $\bm{y}$.
We fit models on several smaller subsets of the ENB data. We randomly select 8, 16, 32, 64, 128, or 256 points as main modality training data ($n_{main}$), with 1, 2, or 4 times as much auxiliary modality data ($r_{aux}$). Auxiliary data are as with other datasets, selected at a superset of input values as the main modality. (Therefore datasets with a ratio of 1:1 auxiliary to main modality data will have all training data for the same input values.) For every combination of these values, we generate 10 datasets by selecting different points from the overall data. For each dataset, we fit a Single, Joint, and Layered model (described in Section 4.2), and record the mean prediction, prediction error, prediction variance, and standardized prediction error for the main modality at each point.
Figure 6: Comparison of average absolute prediction error (left) and relative prediction error (right) for the ENB dataset, by dataset parameters and test category. The “Main” category includes all points where there were main modality (heating load) training data, the “Aux” category includes all points where there were not main modality training data but there were auxiliary modality data, and the “Out of Sample” category includes all other points. There is no “Aux” category where Aux Ratio equals 1, because there were no input values for which only auxiliary modalities were observed in the training data. Additionally, the “Out of Sample” category is missing where the training data, either main or auxiliary modality, encompassed the entire dataset (e.g., in the bottom-right facet, $256\times 4>768$, so no points were fully left out of the training data).
The results of these simulations are summarized in Figure 6. We evaluate prediction accuracy on three subcategories of the dataset: Main, Aux, and Out of Sample. The Main category contains in-sample points where the main modality response was observed in the training data. The Aux category contains in-sample points where the main modality response was not observed in the training data, but auxiliary modalities were. Finally, the Out of Sample category contains all other points where no training data were present. The absolute error and standardized error are averaged over every point in each category, and the figure shows the distribution of this category average over the 10 unique simulations for each combination of $n_{main}$, $r_{aux}$, and model type. We are interested primarily in the multi-modal models’ ability to make improved predictions at points outside of the main modality training data, either with auxiliary training data present (Aux) or without (Out of Sample). A lower absolute prediction error indicates more accurate predictions, on average, and a standardized prediction error closer to 1 indicates properly-calibrated prediction uncertainty. We see improvement in absolute error for multi-modal models over the uni-modal model for larger values of $r_{aux}$. This trend is expected, as the more auxiliary data are present, the more additional information is available for the multi-modal models to inform their predictions. We also see an improvement in the multi-modal models for larger values of $n_{main}$. This trend is also expected, as when $n_{main}$ increases there is a larger amount of paired main and auxiliary training data the model can use to learn relationships between modalities. Overall, both the Joint and Layered models see improved prediction error. However, the Layered model has the largest and most consistent improvements, showing smaller error than both the Single and Joint models for some medium values of $n_{main}$ and $r_{aux}$. From the standardized error, we see overall well-calibrated uncertainty in the Aux and Out of Sample categories (standardized error close to 1). There is a noticeable reduction in standardized error in the multi-modal models for combinations of both large $n_{main}$ and large $r_{aux}$, however, it remains relatively close to 1.
5.2 Jura Dataset
The Jura dataset contains measurements of seven heavy metal concentrations (chromium, nickel, lead, zinc, cadmium, cobalt, and copper) at 359 locations in the Swiss Jura region. In addition to the multiple target variables, each observation contains $X$ and $Y$ geographic coordinates, a categorical variable for one of five rock types, and a categorical variable for one of four land uses. The categorical variables are encoded as five and four indicator variables, respectively. Most observations lie in a geographic grid, with unstructured clusters of additional observations around several points.
This dataset has been analyzed in a multi-target regression context by [24], who identify that the correlations between measurements can be used to predict concentrations of metals that are more expensive to test for using measurements of metals which are less expensive to test for. For our tests, we use the zinc measurement as the scarce main modality and the other metals as the auxiliary modalities. Because each target variable is strictly positive and right-skewed, we log-transform each variable before fitting models. Results are shown for the $\log(\bm{y})$ rather than $\bm{y}$.
We fit models on several smaller subsets of the JURA data, following the same design as with the ENB data. We randomly select 8, 16, 32, 64, 128, or 256 points as main modality training data ($n_{main}$), with 1, 2, or 4 times as much auxiliary modality data ($r_{aux}$). Certain values of rock type and land use are present in very few points in the dataset. In order to avoid imbalanced data problems, and to simplify the demonstration, we exclude the land use and rock type categorical inputs. The imbalanced data issue could also be resolved in other ways such as oversampling and is beyond the scope of this work. For every combination of these values, we generate 10 datasets by selecting different points from the overall data. For each dataset, we fit a Single, Joint, and Layered model (described in Section 4.2), and record the mean prediction, prediction error, prediction variance, and standardized prediction error for the main modality at each point.
Figure 7: Comparison of average absolute prediction error (left) and relative prediction error (right) for the Jura dataset, by dataset parameters and test category. The “Main” category includes all points where there were main modality (Zinc) training data, the “Aux” category includes all points where there were not main modality training data but there were auxiliary modality data, and the “Out of Sample” category includes all other points. There is no “Aux” category where Aux Ratio equals 1, because there were no input values for which only auxiliary modalities were observed in the training data. Additionally, the “Out of Sample” category is missing where the training data, either main or auxiliary modality, encompassed the entire dataset (e.g., in the bottom-right facet, $256\times 4>359$, so no points were fully left out of the training data).
Figure 7 shows an overall summary view of model performance on the Jura data, just as Figure 6 shows on the ENB data. The Main, Aux and Out of Sample categories, and category-wide averages of errors are created in the same manner as with the ENB data. We see some of the same overall trends in average bias as with the ENB data, with multi-modal models showing improvement over uni-modal models for large $n_{main}$ and $r_{aux}$. However, where for the ENB data we saw more comparable improvement between the Joint and Layered models, for the Jura data we see the Layered model outperforming the Joint model in most circumstances. Additionally, the calibration of the uncertainty quantification is generally poorer for the Jura dataset, as shown by the standardized prediction error (right panel) being larger than 1, often between 2 and 3. This indicates that the uncertainty estimate for predictions is generally smaller than it should be if it were properly calibrated. This discrepancy could be due to the fact that the Jura data are measurements, while the ENB data are from simulations. Even though the standardized error is high, it is still closer to 1 for the multi-modal models than for the uni-modal model, especially for larger values of $r_{aux}$.
Figure 8: A comparison of prediction bias (average absolute error) for each multi-modal model vs the uni-modal model on one instance of a trained model with $n_{main}=32$ and $r_{aux}=4$. Positive values (green) mean the multi-modal model reduced average error versus the uni-modal model, while negative values (red) mean the multi-modal model increased average error. The locations of main modality training data are shown as stars, and the locations of auxiliary training data are shown as circles. Points co-located with auxiliary data are predominantly green in the Layered model and green or yellow in the Joint model, indicating that predictions were improved relative to the uni-modal model.
We can take advantage of the Jura dataset’s two-dimensional spatial input values to visualize multi-modal model performance in one trained instance. Figure 8 shows a comparison of main modality prediction bias for one instance of the multi-modal models compared to the uni-modal model. Qualitatively, we see more green points inside the circles, i.e., where auxiliary modality responses were included in training data but the main modality response was not. We also see in some areas between groups of auxiliary observations, the points are also more often green, indicating that the auxiliary training data improves interpolation to out-of-sample test points.
@@ 6 Conclusion
In this paper, we have developed two novel Bayesian neural network-based multi-modal surrogate models. These models build upon existing work in multi-modal learning, Bayesian neural networks, and multi-fidelity surrogate modeling. We estimate the parameters of these models using variational Bayes estimation of these models’ posterior distributions that leverages known and closed-form expressible posterior distributions in the last layer. We developed a novel technique to complete this estimation procedure in the presence of missing data by drawing the missing values from distributions that approximate the posterior predictive distribution of the missing data. Without this step, the missing data would make the last-layer posterior distributions have no known closed form. We demonstrate through simulation that these models show improved estimation for the main modality in some datasets when additional samples of auxiliary modalities are present. The canonical correlation between the auxiliary modality data, collectively, and the main modality or quantity of interest appears to be indicative to determine a priori whether these multi-modal models may lead to improved predictions compared to uni-modal alternatives. While multi-fidelity surrogate models are often hierarchical, these multi-modal models consider alternate data sources equally and enable utilizing them in combination. A method that measures non-linear relationships in order to determine a priori the performance of these multi-modal models is left for future work.
@@ Acknowledgments
This work was authored by the National Laboratory of the Rockies, operated by Alliance for Energy Innovation, LLC, for the U.S. Department of Energy (DOE) under Contract No. DE-AC36-08GO28308. Funding was provided by the Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program through the FASTMath Institute. The views expressed in this article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes.
This research was performed using computational resources sponsored by the Department of Energy’s Office of Critical Minerals and Energy Innovation and located at the National Laboratory of the Rockies.
@@ Appendix A Simulated Datasets
For the “Time Series” dataset, we use $N=200$. We use a PCA-decomposed representation of the full-length $N=200$ noisy time series as the main modality, and all three scalar values as the auxiliary modalities. For “Time Series (1D)”, we instead use Total Distance as the main modality and the PCA-decomposed time series and the other two scalar values as auxiliary modalities. We have one option with a scalar quantity of interest to better replicate downstream tasks such as optimization. Of all the scalar modalities, total distance is most affected by all of the input values. In both datasets, we observe the main modality at the points $(\log\alpha,\log\gamma,\mathrm{atanh}(\delta))\in\{-1.5,-1,-0.5,0.5,1,1.5\}^{3}$ and the auxiliary modalities at the points $(\log\alpha,\log\gamma,\mathrm{artanh}(\delta))\in\{-2,-1.5,-1,-0.5,0,0.5,1,1.5,2\}^{3}$. The number of PCA components is chosen to explain 95% of the variability in the observed time series. Figure 9 shows an example of the original noisy time series data along with the cumulative sum of the seven principal components that account for 95% of the variation in the training data.